jagr2713 (jagr2713):

Trick Question: There are 12 kids in a classroom. 6 kids are wearing socks and 4 are wearing shoes. 3 kids are wearing both. How many are in bare feet?

OpenStudy (mikeymaximum):

2?

jagr2713 (jagr2713):

Nope

OpenStudy (anonymous):

1

jagr2713 (jagr2713):

Nope

Miracrown (miracrown):

5

pooja195 (pooja195):

5 kids

OpenStudy (anonymous):

12

jagr2713 (jagr2713):

Correct @Miracrown

pooja195 (pooja195):

lag >:(

jagr2713 (jagr2713):

D;

OpenStudy (jhannybean):

Can someone explain how they found their answer?

jagr2713 (jagr2713):

^^

OpenStudy (rational):



jagr2713 (jagr2713):

Niceeeeee @rational

jagr2713 (jagr2713):

Here is why As we know 3 kids are wearing both, So only 3 kids are wearing only socks(6 - 3 = 3) and 1 is wearing only shoes(4 - 3 = 1). So, in total 3 + 3 + 1 = 7. Now 12 are there, so 12 - 7 = 5. 5 kids are in bare feet.

pooja195 (pooja195):

3 kids wearing socks 6 - 3 = 3 l w/shoes 4 - 3 = 1 3 + 3 + 1 = 7. 12 are there, so 12 - 7 = 5

OpenStudy (jhannybean):

Good, thank you.

jagr2713 (jagr2713):

GREAT :d